OpenStudy (brucebaner):

The amount of money an electrician charges when visiting a home to make a repair is shown by the linear equation c = 110 + 85h, where c is the amount the electrician charges and h is the number of hours spent at the home. How much does the electrician charge when h = 3.5?

OpenStudy (brucebaner):

c = 110 + 85h c = 110 + 85(3.5) c = 110 + 297.5 c = 407.5

OpenStudy (brucebaner):

@KamiBug

OpenStudy (kamibug):

\(\huge\cal\color{indigo}{Correct!~:)}\)